OpenStudy (anonymous):

tan 2x- cot x=0 in the iinterval 0,2pi

OpenStudy (psymon):

Well no matter what, you have to reduce tan(2x) down to a single angle, not 2 like you have. Happen to have the double angle formula for tangent handy on ya?

OpenStudy (anonymous):

yes, 2 tan u/1-tan^2 u

OpenStudy (psymon):

Alright, awesome. So since doing that identity gives ua a bunch of tangents, I would suggest turning your cotangent into a tangent as well.

OpenStudy (anonymous):

okay, still not seeing a way to finish it tho

OpenStudy (psymon):

We'll get there :P So we can turn cot into tan by saying cotx = 1/tanx, so we would then have this so far \[\frac{ 2tanx }{ 1-\tan^{2}x }- \frac{1}{tanx}\]From here I would combine this into one fraction by multiplying top and bottom of the left fraction by tanx and by multiplying top and bottom of the right fraction by 1-tan^2(x), giving us: \[\frac{ 2tanx(tanx)-1(1-\tan^{2}x) }{ tanx(1-\tan^{2}x) } \implies \frac{ 3\tan^{2}x - 1 }{ tanx(1-\tan^{2}x)}= 0\] Now from here we need to realize that solving the denominator does us no good. Why? Well, if we multiplied both sides by the denominator, it would disappear. If we set the denominator equal to 0, we'd only get undefined values. Basically at this point, our denominator means nothing. So we just want to solve 3tan^2(x) - 1 = 0

OpenStudy (anonymous):

so essentially where does tan equal 1/3?

OpenStudy (psymon):

Where does tan^2(x) = 1/3. We still have the 2nd power there to deal with.

OpenStudy (anonymous):

ya thats what i was thinking lol

OpenStudy (psymon):

\[tanx = \pm \sqrt{\frac{1}{3}}\implies tanx = \pm \frac{ \sqrt{3} }{ 3 }\]

OpenStudy (anonymous):

hmph

OpenStudy (anonymous):

what gave u the clues to do that

OpenStudy (psymon):

To do what exactly?

OpenStudy (anonymous):

i guess to make it into one fraction, but it only makes sense to make it into one fraction.

OpenStudy (anonymous):

cause you are still just subtracting them

OpenStudy (psymon):

Not only that, but in my experience doing that usually just turns out to be the best the best thing to do. Especially since we have a bunch of tangents, you put them into one fraction and see what can simplify. I also know that when I do so, whatever is in the denominator becomes worthless to solving the problem.

OpenStudy (anonymous):

okay makes sense

OpenStudy (anonymous):

okay here is another question, just to rewrite the expression (cos x+sinx)(cos x -sinx) which would factor into sin^2 x -cos^2 x

OpenStudy (anonymous):

rewrite it with double angle formulas

OpenStudy (psymon):

cos^2(x) -sin^(2)x

OpenStudy (psymon):

Which then yeah, goes to simply cos2x

OpenStudy (anonymous):

why did i not see that?

OpenStudy (psymon):

Dunno, lol. But yeah, even looking at the foil, neither cosine is negative, so no way cos^2 would ever be negative.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so if you have to use the power reducing formulas on cos^4 2x could you just do the 1+cos 2x/4?

OpenStudy (psymon):

Nah. Youd have to say this: \[\cos^{4}2x = (\cos^{2}2x)^{2}= (\frac{ 1+\cos2x }{ 2 })^{2}\]

OpenStudy (psymon):

Whoops, correction.

OpenStudy (anonymous):

okay

OpenStudy (psymon):

\[(\frac{ 1+\cos4x }{ 2 })^{2}\]

OpenStudy (anonymous):

thank you so much you have been a GREAT help

OpenStudy (psymon):

Yeah, np ^_^

OpenStudy (anonymous):

Wish u were at my school

OpenStudy (psymon):

They dont have any form of tutorial services at your school?

OpenStudy (anonymous):

im sure they do, but i went to the teacher n thought i had it so i didnt look into it, and i was doing my homework and of course, he didnt explain well enough, so i was looking for all sources possible

OpenStudy (anonymous):

at midnight you know how it is

OpenStudy (psymon):

Yeah. And this is a good place to come to when you cant get any other help. Yeah, either because its midnight or you just want to post a question before you go to bed/school/work and see if someone asnwered it later, anything.

OpenStudy (anonymous):

very good place i love it. i learn more here than i do from teachers.

OpenStudy (psymon):

A lot of place and people who do tutoring or help on the side are better than teachers, sadly enough x_X

OpenStudy (anonymous):

true to a lot of things.

OpenStudy (anonymous):

 what are those points on the x axis y axis they are 1 0 -1 0 1